• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name peter

    Log entry time 08:11:40 on October22,2002

    Entry number 88770

    keyword=shift summary

    01:10 Due to beam trips the average current in the end of run file seems to be lower than expected. We decided to extend the length of each run to try to compensate. Thus, run 2651 is a repeat of 2650.

    01:51 MCC are having problems delivering high current so they are reducing the maximum delivered current to around 85uA - see run 2653.

    02:20 Eric has observed that the target boiling measurements are increasing with beam current - the opposite of what one might expect; see separate HALOG entry for details.

    02:33 MCC telephoned to say the beam would be off for a few minutes as they were having problems with their vacuum.

    02:42 MCC called back to explain that their beam halo was scraping the beamline somewhere and they need the optics group to investigate the problem in the morning. In the meantime, MCC thinks they can deliver about 70uA maximum current; their problems seem to arise at around 80uA.

    03:40 Beam restored by MCC to 20uA. Going to reduce the raster size and compare to previous target boiling run to check rates.

    03:52 Noticed TS1 had crashed and needed rebooted.

    04:12 Finally rebooted TS1 and reset the runcontrol. Going to take a test run at 6uA to check the rates for the target boiling test.

    05:22 Repeating some of the previous target boiling test points again with the 2x2 raster, we STILL have the counts per coulomb INCREASING with beam current; completely the opposite behaviour from what we expect.

    06:45 Mark Jones has suggested taking data with the 15cm dummy target and also to go back to doing target boiling tests with the 15cm liquid hydrogen target again.

    07:01 Take a run with the 15cm dummy target.

    07:30 Move the 15cm liquid hydrogen target into the beam path; liquid deuterium target put to sleep.

    08:01 MCC delivered 110uA beam. Take data on LH2 target with 2x2 raster.

    Run History

    2645 40uA Junk - stopped earlier to change prescalers.
    2646 40uA Target boiling test - good.
    2647 50uA Target boiling test - good.
    2648 65 uA Junk - stopped to change prescalers.
    2649 65uA Target boiling test - good.
    2650 80uA Target boiling test - good.
    2651 80uA Repeat target boiling test - good. (Wanting higher average current for the run at this current.)
    2652 90uA Target boiling test - good.
    2653 85uA Target boiling test - good; lower current due to MCC problems.
    2654 99uA Target boiling test - good for 800k events then beam went off (see comments at 02:33 above).
    2655 Junk Could not start run due to TS1 crash.
    2656 Junk Could not start run due to TS1 crash.
    2657 Junk Could not start run due to TS1 crash.
    2658 6uA Raster check - spot size now ~ 2x2.
    2659 6uA Target boiling test - good.
    2660 9uA Target boiling test - good.
    2661 19.3uA Junk - TS1 crash.
    2662 20.1uA Good run - very strange SPOT plot - see separate HALOG entry.
    2663 29uA Target boiling test - good.
    2664 40uA Target boiling test - good.
    2665 50uA Target boiling test - good.
    2666 60uA Target boiling test - good.
    2667 Junk Incorrect prescale factor
    2668 70uA Target boiling test - good.
    2669 80uA Target boiling test - good.
    2670 30uA Test with 15cm dummy target.
    2671 Junk Incorrect prescale factors.
    2672 100uA Liquid hydrogen target in place; boiling test.
    2673 110uA LH2 target boiling test.